# ATOMIUM: Dust and tracers of binarity in the continua

### v1

@@ Acknowledgements.
TD and NS are supported in part by the Australian Research Council through a Discovery Early Career Researcher Award (DE230100183). TC acknowledges funding from the Research Foundation - Flanders (FWO), grant 1166724N. LD and CL acknowledge support from the KU Leuven C1 excellence grant BRAVE C16/23/009. LD acknowledges support from KU Leuven Methusalem grant SOUL METH/24/012, and the FWO research grants G099720N and G0B3823N. FDC is a Postdoctoral Research Fellow of the Research Foundation - Flanders (FWO), grant 1253223N. IEM acknowledges support from ANID/FONDECYT, grant 11240206 ME acknowledges funding from the FWO research grants G099720N and G0B3823N. TJM acknowledges support from STFC grant no. ST/T000198/1 HSPM thanks the Deutsche Forschungsgemeinschaft (DFG) for support through the collaborative research center SFB 1601 (project ID 500700252) subprojects A4 and Inf. JMCP was supported by STFC grant number ST/T000287/1. DP acknowledges Australian Research Council funding via DP220103767. We also thank the Australia French Association for Research and Innovation (AFRAN) for financially supporting the 5th Phantom users workshop. RS’s contribution to the research described in this publication was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA. RS thanks NASA for financial support via GALEX GO and ADAP awards. LS is an FNRS senior researcher. MVdS acknowledges the Oort Fellowship at Leiden Observatory. OV acknowledges funding from the Research Foundation - Flanders (FWO), grant 1173025N. KTW acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation programme (grant agreement number 883867, project EXWINGS). This work is funded by the French National Research Agency (ANR) project PEPPER (ANR-20-CE31- 0002). We thank the Australian-French Association for Research and Innovation (AFRAN) for financial support. The research leading to these results has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (project UniverScale, grant agreement 951549). This paper makes use of the following ALMA data: ADS/JAO.ALMA#2018.1.00659.L and 2019.1.00187.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. We acknowledge excellent support from the UK ALMA Regional Centre (UK ARC), which is hosted by the Jodrell Bank Centre for Astrophysics (JBCA) at the University of Manchester. The UK ARC Node is supported by STFC Grant ST/P000827/1. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government. This research has made use of the SIMBAD database and the VizieR catalogue access tool, CDS, Strasbourg Astronomical Observatory, France (Wenger et al., 2000; Ochsenbein et al., 2000). This research has made use of the International Variable Star Index (VSX) database, operated at AAVSO, Cambridge, Massachusetts, USA. This research made use of Astropy (Astropy Collaboration et al., 2013, 2018), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007) and NumPy (Harris et al., 2020).References
@@
- Danilovich et al. (2024) Danilovich, T., Malfait, J., Van de Sande, M., et al. 2024, Nature Astronomy, 8, 308
@@
- Greaves & Howarth (2000) Greaves, J. & Howarth, J. J. 2000, Journal of the British Astronomical Association, 110, 131
@@
- Maercker et al. (2012) Maercker, M., Mohamed, S., Vlemmings, W. H. T., et al. 2012, Nature, 490, 232
@@
- Montargès et al. (2025) Montargès, M., Malfait, J., Esseldeurs, M., et al. 2025, A&A, submitted
@@ Appendix A Further details of pulsational periods
Some of the stars in our sample have two pulsation periods listed in Table 1. For SRb variables Gru and V PsA, it is possible that these additional periods (of 128 and 105 days, respectively) are in fact aliasing effects, as they are roughly at frequencies of , where the is the peak frequency in the power spectrum of the survey window (roughly 1 yr). T Mic has a main period of 352 days, and a secondary period of 178 days. Although this period may be from a similar window aliasing effect as mentioned above, the light curve appears to be a superposition of two periods suggesting that two pulsation modes are indeed present. U Del is a case of a semiregular variable with a LSP of 1163 d, with its shorter pulsation period being 120 days (Speil 2006; Cadmus 2024).
@@
The Mira variables R Hya, R Aql and W Aql, and the SRa variable S Pav all have been reported to have long-term period changes. R Aql and R Hya have been reported to have shortening periods (Wood & Zarro 1981; Greaves & Howarth 2000; Joyce et al. 2024), possibly owing to undergoing thermal pulses in the recent past (a few hundred or so years ago). This is discussed further in Sect. 5.3. Despite being classified as SRa, S Pav is a case of a “Mira-like” semiregular variable, as described in Kerschbaum et al. (1996). Its visual amplitude is in the Mira regime (¿ 2.5 mag), and it has periods reported between 380–390 days; a thorough investigation of whether its period is indeed varying may be useful. A careful examination of historical light curve of W Aql (from AAVSO) reveals that it is probably a wandering Mira (Templeton et al. 2005), albeit with some unexplained reductions in peak magnitude in the 1970s and again since 2006.
@@ Appendix B Additional tables and plots
In Tables 5, 6, 7, and 8, we give overviews of the continua observed with the ALMA 12m extended, mid, compact, and combined configurations, and with the ACA in Table 9. The compact continuum images are shown in Fig. 11 and the ACA continuum images are shown in Fig. 12. In Table 10 we give details of the imaging properties of the UD-subtracted images. The UD fits themselves are shown with their corresponding visibility amplitudes in Fig. 13.
@@
AH Sco | 739 | 1.9 | |||||
KW Sgr | 382 | 2.7 | |||||
VX Sgr | 1089 | 4.0 |
Star | Beam size | Beam PA | Observing date | MRSa | rms | Peak flux | Dynamic | Spectral |
[″] | [∘] | [yyyy-mm-dd] | [″] | [Jy/beam] | [Jy/beam] | range | index | |
GY Aql | 2019-11-07 | 69 | 7 | |||||
R Aql | 2019-11-19 | 77 | 14 | |||||
IRC10529 | 2019-11-02 | 66 | 5 | |||||
Gru | 2019-10-16 | 64 | 21 | |||||
U Her | 2019-12-12 | 73 | 19 | |||||
T Mic | 2019-11-11 | 66 | 50 | |||||
@@
Star | Beam size | Beam PA | rms | Peak flux | Dynamic |
[mas] | [∘] | [Jy/beam] | [Jy/beam] | range | |
GY Aql | -52.2 | 19.6 | 730 | 37 | |
R Aql | 10.7 | 9.3 | 509 | 55 | |
IRC10529 | -84.5 | 20.1 | 433 | 22 | |
W Aql | 52.2 | 7.2 | 314 | 44 | |
SV Aqr | 63.2 | 10.3 | 112 | 11 | |
U Del | -0.7 | 11.2 | 49 | 4 | |
Gru | 28.6 | 11.7 | 3288 | 280 | |
U Her | -6.8 | 15.5 | 434 | 28 | |
R Hya | 84.1 | 10.7 | 512 | 48 | |
T Mic | 72.8 | 12.2 | 601 | 49 | |
S Pav | 0.6 | 10.9 | 673 | 62 | |
IRC+10011 | 29.9 | 20.3 | 517 | 25 | |
V PsA | 79.7 | 9.7 | 195 | 20 | |
RW Sco | -33.8 | 21.1 | 1243 | 59 | |
KW Sgr | 38.5 | 8.5 | 48 | 6 | |
VX Sgr | 84.4 | 14.9 | 569 | 38 | |
AH Sco | -56.3 | 11.3 | 505 | 45 |
@@ Appendix D ACA Band 3 molecular lines
In addition to the continuum images, we also obtained some line observations using the ACA. We detected lines of 13CO, C17O, 12CS, 32SO, 34SO, and, tentatively, C18O towards at least one source each. For all sources at least one line was detected, although for Gru we only tentatively detect one 13CO line. The molecular line detections, including transition details and integrated fluxes, are given in Table 11. The line spectra are plotted in Figs. 14–19. Table 11 lists the lines detected in our ACA observations and their integrated fluxes.
@@ D.1 Oxygen isotopic ratios
Although the 13CO line has a high SNR (65), the SNR are much lower for C17O (5) and C18O (1.5), when considering spectra extracted for a circular aperture of 10″. This makes the line flux determinations less reliable for the rarer isotopologues, especially C18O, even when we choose the integration limits based on the well-defined limits of the 13CO line ( to km s-1). We find Jy km s-1 and Jy km s-1 with uncertainties based on the rmses of the spectra. This gives a ratio of owing to the large uncertainty on the C18O line. This does not allow us to put tight constraints on the initial mass of IRC10529, and we find , assuming solar metallicity. The fact that we do weakly detect the C18O indicates that the star has not undergone hot bottom burning, which destroys 18O, and hence puts an upper limit on the initial mass of (Karakas & Lugaro 2016).
@@ Appendix E Calculating brightness temperatures
Appendix F Hydrodynamical model

### v2

@@ Acknowledgements.
We thank the anonymous referee for their careful reading and thoughtful comments on the manuscript. TD and NS are supported in part by the Australian Research Council through a Discovery Early Career Researcher Award (DE230100183). TC acknowledges funding from the Research Foundation - Flanders (FWO), grant 1166724N. LD and CL acknowledge support from the KU Leuven C1 excellence grant BRAVE C16/23/009. LD acknowledges support from KU Leuven Methusalem grant SOUL METH/24/012, and the FWO research grants G099720N and G0B3823N. FDC is a Postdoctoral Research Fellow of the Research Foundation - Flanders (FWO), grant 1253223N. IEM acknowledges support from ANID/FONDECYT, grant 11240206 ME acknowledges funding from the FWO research grants G099720N and G0B3823N. TJM acknowledges support from STFC grant no. ST/T000198/1 HSPM thanks the Deutsche Forschungsgemeinschaft (DFG) for support through the collaborative research center SFB 1601 (project ID 500700252) subprojects A4 and Inf. JMCP was supported by STFC grant number ST/T000287/1. DP acknowledges Australian Research Council funding via DP220103767. We also thank the Australia French Association for Research and Innovation (AFRAN) for financially supporting the 5th Phantom users workshop. RS’s contribution to the research described in this publication was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA. RS thanks NASA for financial support via GALEX GO and ADAP awards. LS is an FNRS senior researcher. MVdS acknowledges the Oort Fellowship at Leiden Observatory. OV acknowledges funding from the Research Foundation - Flanders (FWO), grant 1173025N. KTW acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation programme (grant agreement number 883867, project EXWINGS). This work is funded by the French National Research Agency (ANR) project PEPPER (ANR-20-CE31- 0002). We thank the Australian-French Association for Research and Innovation (AFRAN) for financial support. The research leading to these results has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (project UniverScale, grant agreement 951549). This paper makes use of the following ALMA data: ADS/JAO.ALMA#2018.1.00659.L and 2019.1.00187.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. We acknowledge excellent support from the UK ALMA Regional Centre (UK ARC), which is hosted by the Jodrell Bank Centre for Astrophysics (JBCA) at the University of Manchester. The UK ARC Node is supported by STFC Grant ST/P000827/1. Part of this work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government. This research has made use of the SIMBAD database and the VizieR catalogue access tool, CDS, Strasbourg Astronomical Observatory, France (Wenger et al., 2000; Ochsenbein et al., 2000). This research has made use of the International Variable Star Index (VSX) database, operated at AAVSO, Cambridge, Massachusetts, USA. This research made use of Astropy (Astropy Collaboration et al., 2013, 2018), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007) and NumPy (Harris et al., 2020).References
@@
- Bedding et al. (1998) Bedding, T. R., Zijlstra, A. A., Jones, A., & Foster, G. 1998, MNRAS, 301, 1073
@@
- Danilovich et al. (2024) Danilovich, T., Malfait, J., Van de Sande, M., et al. 2024, Nat. Astron., 8, 308
@@
- Esseldeurs et al. (2025) Esseldeurs, M., Decin, L., & ATOMIUM Consortium. 2025, Nature Astronomy, in press
@@
- Greaves & Howarth (2000) Greaves, J. & Howarth, J. J. 2000, JBAA, 110, 131
@@
- Maercker et al. (2012) Maercker, M., Mohamed, S., Vlemmings, W. H. T., et al. 2012, Nat. Astron., 490, 232
@@
- Montargès et al. (2025) Montargès, M., Malfait, J., Esseldeurs, M., et al. 2025, A&A, 699, A22
@@ Appendix A Further details of pulsational periods
A.1 Overview of pulsational periods
Here we clarify some details for stars which have two or more pulsation periods reported in the literature, and stars which are marked with an asterisk (*) in Table 1. For the SRb variables Gru and V PsA, it is possible that previously reported additional periods (of 128 and 105 days, respectively) are in fact aliasing effects, as they are roughly at frequencies of , where the is the peak frequency in the power spectrum of the survey window (roughly 1 yr). T Mic has a main period of 352 days, and a secondary period of 178 days. Although this period may be from a similar window aliasing effect as mentioned above, the light curve appears to be a superposition of two periods suggesting that two pulsation modes are indeed present. U Del is a case of a semiregular variable with a LSP of 1163 d, with its shorter pulsation period being 120 days (Speil 2006; Cadmus 2024).
@@ A.1 Overview of pulsational periods
The Mira variables R Hya, R Aql and W Aql, and the SRa variable S Pav all have been reported to have long-term period changes. R Aql and R Hya have been reported to have shortening periods (Wood & Zarro 1981; Greaves & Howarth 2000; Joyce et al. 2024), possibly owing to undergoing thermal pulses in the recent past (a few hundred or so years ago). This is discussed further in Sect. A.2. Despite being classified as SRa, S Pav is a case of a “Mira-like” semiregular variable, as described in Kerschbaum et al. (1996). Its visual amplitude is in the Mira regime (¿ 2.5 mag), and it has periods reported between 380–390 days; a thorough investigation of whether its period is indeed varying may be useful. A careful examination of historical light curve of W Aql (from AAVSO) reveals that it is probably a wandering Mira (Templeton et al. 2005), albeit with some unexplained reductions in peak magnitude in the 1970s and again since 2006.
@@
A.2 R Aql and R Hya: recent thermal pulses?
If R Aql and R Hya both underwent a relatively recent (past few hundred years) thermal pulse, as suggested by Wood & Zarro (1981), Greaves & Howarth (2000) and Zijlstra et al. (2002), then we would perhaps expect to see some similarities in their circumstellar envelopes. Their continuum images are indeed qualitatively similar. Although R Hya (Fig. LABEL:fig:rhya) has a larger continuous region of extended flux than R Aql (Fig. 2), both stars have some clumpy regions of continuum flux offset from the continuum peak (e.g. up to ), most likely originating from dust clumps. The ALMA continuum observations are in general agreement with the SPHERE observations of Montargès et al. (2023), which reveal a larger region of polarised flux for R Hya than for R Aql.
The physical radii of both stars, as derived from our UD fits, and their monochromatic luminosities are close to being the same, within our uncertainties. They are also the two Mira stars with the smallest radii when comparing physical units, both from our UD fits (Table 2) and from literature measurements taken in the NIR (Table 4). From their study of dust in the infrared, Zhao-Geisler et al. (2012) estimated that the inner edges of the IR dust shells around these stars are au and au from the corresponding stars, for R Aql and R Hya respectively. Strikingly, R Aql shows an increase in flux in the far infrared (in the ISO LWS spectrum, see Figure 2 of Zhao-Geisler et al. 2012), suggesting a shell of cool dust, as might be expected for a detached (or detaching) dust shell. The same is not seen for R Hya (see also the earlier results of Decin et al. 2008). If R Aql truly has a burgeoning detached shell while R Hya — which has a significant amount of dust close to the star — does not, that would imply that R Hya had a more recent thermal pulse. This is counter to the analysis based on luminosity done by Wood & Zarro (1981), which predicts that less time has passed since the thermal pulse of R Aql than R Hya. However, Maercker et al. (2016, 2024) presented ALMA observations that demonstrate that for some of the surveyed (carbon-rich) AGB stars the shells thought to have been formed through thermal pulses (1000–8000 years ago) were followed by a moderate decrease in mass loss, rather than an abrupt decrease seen in other cases. It is possible that R Hya has undergone a more moderate decrease in mass-loss post-thermal pulse, while R Aql has undergone a more drastic change. The mass-loss rates calculated by Bergman & Humphreys (2020) show a mass-loss rate for R Aql about five times higher than for R Hya (see Table 14), which could be because the historic mass loss probed by CO lines mainly captures the imprint of enhanced mass loss during the thermal pulse for R Aql, but more of the lower non-thermal pulse generated mass loss for R Hya.
On larger scales than our ALMA observations, a dust shell around R Hya was proposed by Hashimoto & Izumiura (1998) based on IRAS spectroscopy and photometric imaging. Higher resolution Herschel/PACS imaging reveals a fermata-style bow shock around R Hya (Cox et al. 2012), which Maercker et al. (2022) modelled as a partial ellipsoid shell approximately 100″from the AGB star. This more extended dust cannot be captured in our ALMA observations, partly because much of it falls outside of our field of view, and partly because large scale structures tend to be resolved out by ALMA. Similar observations of R Aql are not available; R Aql was not imaged with Herschel/PACS.
The CO channel maps of of R Aql and R Hya, presented in Decin et al. (2020), show unique patterns for both stars, dissimilar to each other and to the other stars in the ATOMIUM sample. R Aql has a rose-like pattern, which Malfait et al. (2024b) was unable to reproduce through a hydrodynamical model of a hierarchical triple star system161616Malfait et al. (2024a) found that a qualitatively similar pattern seen in the earlier hydrodynamic models of El Mellah et al. (2020) and Malfait et al. (2021) was most likely caused by periodic instabilities arising from an absence of cooling functions aside from adiabatic expansion.. A possibility is that the rose-like structure was caused by a thermal pulse and that the irregular arc-like pattern is an imprint of the convective flows in the stellar atmosphere (compare with the gas densities in the models of Freytag & Höfner 2023 and the irregularities in the otherwise thin detached shells seen by Maercker et al. 2024). From the angular extent of the shapes in the CO emission, the distance to the star, and the expansion velocities (Decin et al. 2020; Wallström et al. 2024), this would imply a thermal pulse –1000 years ago.
R Hya has at least three pairs of bipolar structures, as described in detail by Homan et al. (2021). Decin et al. (2008) estimated a recent decrease in mass loss by a factor of 20 for R Hya, which they attribute to a thermal pulse period ending 220 years ago, in agreement with the timeline proposed by Wood & Zarro (1981). However, it is unclear how the complicated circumstellar structure revealed by the ATOMIUM molecular lines (Homan et al. 2021) and the dust in the continuum observations (Fig. LABEL:fig:rhya) can be reconciled with a recent thermal pulse. A possible explanation could be a companion star in the inner wind, disrupting the more spherical shell that is expected to form during a thermal pulse.
@@ Appendix B Additional tables and plots
In Tables 7, 8, 9, and 10, we give overviews of the continua observed with the ALMA 12m extended, mid, compact, and combined configurations, and with the ACA in Table 11. The compact continuum images are shown in Fig. 14 and the ACA continuum images are shown in Fig. 15. In Table 12 we give details of the imaging properties of the UD-subtracted images. The UD fits themselves are shown with their corresponding visibility amplitudes in Fig. 16.
In Fig. 17 we reproduce Fig. 7 with the addition of the literature data given in Table 6. See Sect. 5.2 for details.
@@
| KW Sgr | 382 | 0.6 | |||||
| VX Sgr | 1089 | 4.0 | |||||
| AH Sco | 739 | 1.9 |
| Star | Beam size | Beam PA | Observing date | MRSa | rms | Peak flux | Dynamic |
| [″] | [∘] | [yyyy-mm-dd] | [″] | [Jy/beam] | [Jy/beam] | range | |
| GY Aql | 2019-11-07 | 69 | 7 | ||||
| R Aql | 2019-11-19 | 77 | 14 | ||||
| IRC10529 | 2019-11-02 | 66 | 5 | ||||
| Gru | 2019-10-16 | 64 | 21 | ||||
| U Her | 2019-12-12 | 73 | 19 | ||||
| T Mic | 2019-11-11 | 66 | 50 | ||||
@@
| Star | Beam size | Beam PA | rms | Peak flux | Total flux | Dynamic |
| [mas] | [∘] | [Jy/beam] | [Jy/beam] | [mJy] | range | |
| GY Aql | 19.6 | 730 | 2.90 | 37 | ||
| R Aql | 11.6 | 437 | 9.60 | 38 | ||
| IRC10529 | 20.1 | 433 | 4.93 | 22 | ||
| W Aql | 7.2 | 314 | 4.96 | 44 | ||
| SV Aqr | 10.3 | 112 | 0.47 | 11 | ||
| U Del | 11.2 | 49 | 2.31 | 4 | ||
| Gru | 11.7 | 3288 | 7.06 | 280 | ||
| U Her | 15.5 | 434 | 6.18 | 28 | ||
| R Hya | 10.7 | 512 | 6.93 | 48 | ||
| T Mic | 12.2 | 601 | 3.53 | 49 | ||
| S Pav | 10.9 | 673 | 5.34 | 62 | ||
| IRC+10011 | 20.3 | 517 | 3.07 | 25 | ||
| V PsA | 9.7 | 195 | 1.62 | 20 | ||
| RW Sco | 21.1 | 1243 | 0.77 | 59 | ||
| KW Sgr | 8.5 | 48 | 0.61 | 67 | ||
| VX Sgr | 14.9 | 569 | 3.87 | 34 | ||
| AH Sco | 11.3 | 505 | 5.39 | 45 |
@@ Appendix D ACA Band 3 molecular lines
In addition to the continuum images, we also obtained some line observations using the ACA. We detected lines of 13CO, C17O, 12CS, 32SO, 34SO, and, tentatively, C18O towards at least one source each. For all sources at least one line was detected, although for Gru we only tentatively detect one 13CO line. The molecular line detections, including transition details and integrated fluxes, are given in Table 13. The line spectra are plotted in Figs. 20–25. Table 13 lists the lines detected in our ACA observations and their integrated fluxes.
@@ D.1 Oxygen isotopic ratios
Although the 13CO line has a high S/N (65), the S/N are much lower for C17O (5) and C18O (1.5), when considering spectra extracted for a circular aperture of 10″. This makes the line flux determinations less reliable for the rarer isotopologues, especially C18O, even when we choose the integration limits based on the well-defined limits of the 13CO line ( to km s-1). We find Jy km s-1 and Jy km s-1 with uncertainties based on the rmses of the spectra. This gives a ratio of owing to the large uncertainty on the C18O line. This does not allow us to put tight constraints on the initial mass of IRC10529, and we find , assuming solar metallicity. The fact that we do weakly detect the C18O indicates that the star has not undergone hot bottom burning, which destroys 18O, and hence puts an upper limit on the initial mass of (Karakas & Lugaro 2016).
@@ Appendix E Calculating brightness temperatures
Appendix F Apparent extent of bright unresolved sources
The apparent size of a source observed with ALMA (or another radio interferometer) depends on the true size of the source, the dimensions of the synthetic beam, and the intensity of the source. We can think of the latter quantity in terms of the dynamic range or the S/N ratio. When giving the size of the synthetic beam (as in Tables 7–11) we give the FWHM along the major and minor axes. For very bright sources, the apparent size of the emission when considering the or contours, where is the rms noise, can be much larger than the beam, even if the source itself is smaller than the beam. If we consider a Gaussian distribution of flux along one of the beam axes:
where is the peak flux, is the offset from the centre of the beam, is the FWHM of the beam along the major (or minor) axis, and , chosen to account for the fact that is related to the Gaussian standard deviation, , by . We can then calculate that, for a point source, the contour, where is the rms noise, will lie at an offset that depends on and the dynamic range of the observation, . In this case, and we are interested in . Then equation 12 can be rearranged to find
From this we can see that for a very bright point source (e.g. large ), the contour would be located at an offset of from the centre of the emission. Such a source would appear to have a diameter of . Thus, we cannot conclude that such a source is resolved if the contour has a diameter of . This effect can be significant even for more modest dynamic ranges; for example, for , the contour would have a diameter of around .
Appendix G Mass-loss rates
| Star | Mass-loss rate | Reference |
| GY Aql | Wannier & Sahai (1986) | |
| R Aql | Bergman & Humphreys (2020) | |
| IRC10529 | Danilovich et al. (2015b) | |
| W Aql | Ramstedt et al. (2017) | |
| SV Aqr | Olofsson et al. (2002) | |
| U Del | Olofsson et al. (2002) | |
| Gru | Doan et al. (2017) | |
| U Her | Bergman & Humphreys (2020) | |
| R Hya | Bergman & Humphreys (2020) | |
| T Mic | Olofsson et al. (2002) | |
| S Pav | Olofsson et al. (2002) | |
| IRC+10011 | Danilovich et al. (2017b) | |
| V PsA | Olofsson et al. (2002) | |
| RW Sco | Groenewegen et al. (1999) |
G.1 Notes on a few individual SUCCESS stars in light of binary-enhanced dust formation
An interesting example is the dusty but relatively low mass-loss rate AGB star L2 Pup (the lowest mass-loss rate source in Fig. 26). Kervella et al. (2016) found a low mass of for the companion, which is embedded in a dusty, nearly edge-on disc (Kervella et al. 2015; Ohnaka et al. 2015). The dust around L2 Pup is not distributed spherically, suggesting that the companion has strongly influenced its distribution. However, with a low mass, the gravitational potential of the companion is low, and it will cause weaker shocks as it orbits the AGB star (Maes et al. 2021), contributing to less dust formation than if it were more massive. This explains why the mass-loss rate of L2 Pup remains low (, Danilovich et al. 2015b and Appendix G).
We also note that the carbon star with the highest mass-loss rate in the SUCCESS sample, comparable to the aforementioned oxygen-rich stars, is II Lup, around which Unnikrishnan et al. (2024) recently tentatively identified SiN, which suggests the presence of a binary companion (Van de Sande & Millar 2022). The oxygen-rich star with the highest mass-loss rate in the SUCCESS sample is GX Mon, around which earlier observations have shown a circumstellar spiral, also indicative of a binary companion (Randall et al. 2020).
Appendix H Hydrodynamical model